OpenStudy (anonymous):

@amistre64 Can you please explain the Rational Roots Theorem? Example: List the possible rational roots of 3x^3+5x^2-2x+1=0

OpenStudy (amistre64):

yay! no spinning wheel of death ...

OpenStudy (anonymous):

Lol?

OpenStudy (amistre64):

do you recall a method named "FOIL"?

OpenStudy (anonymous):

YES :)

OpenStudy (amistre64):

the key letters in that is the F and L ... at least for this purpose

OpenStudy (amistre64):

assume we have some factors:\[(ax-b)(cx-d)\]or expanded to be:\[(ac)x^2+(-ad-bc)x+(bd)\] have you gone over the quadratic formula yet?

OpenStudy (anonymous):

Yes I have.

OpenStudy (amistre64):

if we run that thru the quadratic formula; we would get zeros that are formed from the factors of the last term, \(bd\), divided by the factors of the first term \(ac\). This works for any polynomial in general.

OpenStudy (amistre64):

so if there is a rational solution to this setup, it would have to be a product of:\[\frac{factors~of~last~term}{factors~of~first~term}\]

OpenStudy (anonymous):

Okay, so I run the equation through the quadratic formula and... My solution is supposed to look like (factors of last term)/(factors of first term)?

OpenStudy (amistre64):

the quadratic only works for degree 2 polynomials ... but its a way to help visualise the rational roots thought process

OpenStudy (amistre64):

3x^3+5x^2-2x+1=0 last term is 1; factors of 1 are: 1 first term is 3; factors of 3 are: 1,3 if there is a rational root to this setup, it will have to be from a pool of options:\[\pm\frac{1}{1},\pm\frac{1}{3}\]

OpenStudy (anonymous):

Oooooh. That's not as complicated as I thought...

OpenStudy (amistre64):

yeah, it is a pretty simple idea :)

OpenStudy (anonymous):

So for x^3+2x+3=0... I do the same thing? First and last factors?

OpenStudy (amistre64):

yep

OpenStudy (anonymous):

So I have the same answer then.

OpenStudy (anonymous):

Huh-uh... No. Sorry.

OpenStudy (amistre64):

not quite. last/first factors of last are 1,3 factors of first are 1 so the possible solutions are: +- 1, +- 3

OpenStudy (amistre64):

the rational roots creates a set of possible solutions, but it does not guarentee that any of them are actually a solution. IF it has a rational solution, then it would have to be one of the options in the pool; but there is no certainty that there is a rational solution to begin with.